The perimeter of a square floor is 68 feet. What is the areas of the floor? Answer in units of ft2.
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

Perimeter of a square is 4s, where s is side length
68/4=17 feet per side
Area is s^2, so that is 17^2=289 sq ft.

Answer by Alan3354(69443)

The perimeter of a square floor is 68 feet. What is the areas of the floor? Answer in units of ft2.
============
A square has 4 sides of equal length.
Find the length of the sides.
Call it s.
==============
Area = s*s = s^2